# Strong coupling non-Markovian quantum thermodynamics using exact dynamics of a central spin surrounded by a spin bath and the role of a canonical Hamiltonian

### v1

@@ Abstract
The focus is on understanding the quantum thermodynamics of strongly coupled non-Markovian quantum systems. To this end, a non-trivial, non-Markovian model of a central spin surrounded by a spin bath is taken up, and its exact evolution is derived for arbitrary system-bath couplings. The fundamental quantum thermodynamic quantities, such as system and bath internal energies, work, heat, entropy production, and ergotropy, are calculated using the dynamics and original system (bath) Hamiltonian. An explicit expression for the work, a mismatch between the system and bath internal energies, is derived. Further, an interesting observation relevant to the spin bath acting as a charger is made in a scenario where the central spin is envisaged as a quantum battery. The role of a canonical Hamiltonian in calculating the above thermodynamic quantities, a recently developed technique, is also investigated.
@@ I Introduction
One of the cornerstones of the field of strong-coupling non-Markovian quantum thermodynamics is to correctly define the fundamental thermodynamic quantities such as heat, work, internal energy, ergotropy, and entropy production in the quantum regime [62, 63, 64, 65, 66]. This is essential for the definition of thermodynamic laws of strongly coupled NM quantum systems. In this context, a recently developed technique of canonical Hamiltonian and minimal dissipator constructed with traceless Lindblad operators [67] has been utilized to address these issues [68, 69]. Here, the canonical Hamiltonian rather than the system self-Hamiltonian is used to study the impact of strong coupling and non-Markovian effects on the system. When in a strong coupling situation, the system evolves under the modified ‘canonical Hamiltonian,’ which contains information about its energies.
@@
In this work, we study the strong-coupling non-Markovian quantum thermodynamics of a central spin interacting with a spin bath [70, 71, 72, 73, 74]. The baths can be broadly placed into two categories [75]: (a) Bosonic and (b) spin baths. The spin-boson model [8] and the Caldeira-Leggett model [11] are well-known examples of bosonic baths. The dynamics of these types of models have been studied in the literature [1, 2, 3]. Conversely, when dealing with spin baths [76], frequent use is made of perturbative methods or time-dependent quantum master equations [77, 78, 79]. Studying these systems is crucial for understanding magnetic systems [80], quantum spin glasses [81], quantum batteries [82], NV-center [83], and superconducting systems [77], among others.
@@
Here, we develop the exact reduced dynamics of the central spin model. This is then used to understand the quantum thermodynamic properties of the central spin. To this effect, we calculate the thermodynamic quantities of this system using two methods. First, a use is made of the system and the bath Hamiltonian, where the work done by the bath can be seen as the mismatch between the change in the system and the bath’s internal energies. This is next benchmarked by adopting the canonical Hamiltonian, as the replacement for the system’s Hamiltonian [68], to calculate the corresponding quantities. The aim is to arrive at the correct forms of the quantities, such as heat, work, internal energy, entropy production, and ergotropy in strongly coupled non-Markovian quantum systems and the correct forms of the corresponding laws of quantum thermodynamics, taking the central spin model as an example.
The work is organized as follows. In Sec. II, we briefly discuss the theory of canonical Hamiltonian. The exact dynamics of the central spin model and the corresponding canonical Hamiltonian are developed in Sec. III. Section IV discusses the quantum thermodynamic quantities of the system using the original system and the bath Hamiltonian, where we calculate the system’s heat, work, internal energy, entropy production, and ergotropy. In Sec. V, the same calculations are revisited using the canonical Hamiltonian approach. This is followed by the conclusion in Sec. VI.
II The theory of canonical Hamiltonian
@@
$H^{\rm canonical}$ $=\frac{1}{2id}\sum_{j}\gamma_{j}\bigg{(}{\rm Tr}(E_{j})E_{j}^{\dagger}-{\rm Tr}(E_{j}^{\dagger})E_{j}\bigg{)},$ (7)
@@
We shall next calculate the canonical Hamiltonian of a central spin surrounded by a spin bath, where the central spin interacts uniformly with the bath.
III The central spin model
@@
where $\omega_{0}$ is the transition frequency of the central spin-$1/2$ particle, $\epsilon$ is the interaction strength between the system and the bath with $\sqrt{N}$ being the scaling factor. $\omega/N$ is the scaled frequency of the bath. The evolution of the composite system $S$-$B$ for an arbitrary initial state $\rho_{SB}(0)=\rho_{S}(0)\otimes\rho_{B}(0)$, assuming separable initial condition, of the total system (where $\rho_{S}(0)$ and $\rho_{B}(0)$ are the initial states of the central spin and the bath, respectively) by means of the global unitary $U=e^{-iHt}$ is given by
@@
where $H$ is the total Hamiltonian. Here, we consider the Gibbs state as the initial state of the bath, which is given by
@@
$S_{n}(t)$ $=\sum^{N}_{n=1}e^{-\frac{\beta\omega}{2}\left(1-\frac{2n}{N}\right)}\left\{\left(\frac{\chi_{+}(n)^{2}}{1+\chi_{+}(n)^{2}}\right)^{2}+\left(\frac{\chi_{-}(n)^{2}}{1+\chi_{-}(n)^{2}}\right)^{2}+\left(\frac{\chi_{+}(n)^{2}\chi_{-}(n)^{2}}{\left[1+\chi_{+}(n)^{2}\right]\left[1+\chi_{-}(n)^{2}\right]}\right)2\cos\left(\frac{t\sqrt{b_{n}^{2}+4a_{n}^{2}}}{N}\right)\right\}.$ (16)
@@
This is used to find out the canonical Hamiltonian of the central spin model.
III.1 The canonical Hamiltonian

### v2

# Strong coupling non-Markovian quantum thermodynamics of a finite-bath system

@@ Abstract
The focus is on understanding the quantum thermodynamics of strongly coupled non-Markovian quantum systems. To this end, a non-trivial, non-Markovian model of a central spin surrounded by a spin bath is taken up, and its exact evolution is derived for arbitrary system-bath couplings. The fundamental quantum thermodynamic quantities, such as system and bath internal energies, work, heat, entropy production, and ergotropy, are calculated using the dynamics and original system (bath) Hamiltonian. An explicit expression for the work, a mismatch between the system and bath internal energies, is derived. The thermodynamic entropy of the system at thermal equilibrium is studied using the Hamiltonian of mean force in the strong coupling regime. The role of a canonical Hamiltonian in calculating the above thermodynamic quantities, a recently developed technique, is also investigated. Further, an interesting observation relevant to the spin bath acting as a charger is made in a scenario where the central spin is envisaged as a quantum battery.
@@ I Introduction
One of the cornerstones of the field of strong-coupling non-Markovian quantum thermodynamics is to correctly define the fundamental thermodynamic quantities such as heat, work, internal energy, ergotropy, and entropy production in the quantum regime [62, 63, 64, 65, 66, 67]. This is essential for the definition of thermodynamic laws of strongly coupled NM quantum systems. Further, at strong system and bath couplings, the equilibrium state of the system coupled to a bath has been found to be set by the Hamiltonian of mean force (HMF) instead of the bare Hamiltonian of the system [68, 31, 69]. Moreover, a recently developed technique of canonical Hamiltonian and minimal dissipator constructed with traceless Lindblad operators [70] has been utilized to calculate the thermodynamic quantities of the system [71, 72], where the canonical Hamiltonian rather than the system’s bare Hamiltonian is used to study the impact of strong coupling and non-Markovian effects on the system.
@@
In this work, we study the strong-coupling non-Markovian quantum thermodynamics of a central spin interacting with a spin bath [73, 74, 75, 76, 77]. The baths can be broadly placed into the following categories [78]: (a) Bosonic, (b) fermionic, and (c) spin baths. The spin-boson model [8] and the Caldeira-Leggett model [11] are well-known examples of bosonic baths. The dynamics of these types of models have been studied in the literature [1, 2, 3]. Conducting electrons in the Anderson and the Kondo models are examples of a fermionic bath system. Numerically exact methods, for example, the fermionic hierarchical equation of motion, have been used to obtain the dynamics of the above systems [79, 80]. Further, exact dynamics of noninteracting fermions (bosons) linearly coupled to thermal environments of noninteracting fermions can be found in [81]. Conversely, when dealing with spin baths [82], frequent use is made of perturbative methods or time-dependent quantum master equations [83, 84, 85]. Dissipative equation of motion approach has also been used to find exact dynamics of the systems surrounded by spin bath [86]. Studying these systems is crucial for understanding magnetic systems [87], quantum spin glasses [88], quantum batteries [89], NV-center [90], and superconducting systems [83], among others. Moreover, the finite nature of the bath renders the exact dynamics of the central spin model analytically tractable across arbitrary system-bath coupling regimes. This unique feature provides profound advantages for probing the non-Markovian effects on quantum thermodynamics, particularly in strong coupling regimes.
@@
Here, we develop the exact dynamics of the central spin model. This is then used to understand the quantum thermodynamic properties of the central spin. To this effect, we calculate the thermodynamic quantities of this system using a number of methods outlined above. We begin by analyzing entropy production, heat, and work while discussing the first and second laws of thermodynamics. We also find out the cost of turning the system and bath interactions on or off, and we explore a variation of the second law in finite bath models. Next, we calculate the thermal equilibrium state and entropy under strong coupling using the HMF technique, which aids in examining the third law of quantum thermodynamics. We then discuss the use of the canonical Hamiltonian as the replacement for the system’s bare Hamiltonian to calculate the thermodynamic quantities. We aim to arrive at the correct forms of the quantities, such as heat, work, internal energy, entropy production, and thermal equilibrium state in strongly coupled non-Markovian quantum systems and the correct forms of the corresponding laws of quantum thermodynamics, taking the central spin model as an example. Finally, we demonstrate the central spin model as a quantum battery, highlighting how a finite spin bath uniquely charges the central spin quantum battery. To summarize, we first present a brief discussion of the tools for handling strong-coupling non-Markovian quantum thermodynamics (Sec. II). This is followed in the later sections with new results on the exact dynamics of the central spin model and its quantum thermodynamics in the strong coupling non-Markovian regime, by applying the above tools.
The organization of this work is as follows: In Sec. II, we discuss tools for exploring non-Markovian strong-coupling quantum thermodynamics. This includes discussions on entropy production, heat, work, the Hamiltonian of mean force, and the canonical Hamiltonian technique. Section III covers the exact dynamics of the central spin model. In Section IV, we analyze thermodynamic quantities in the strong coupling and non-Markovian regime of the central spin model. The application of the central spin model as a quantum battery, along with the calculation of its ergotropy, is presented in Section V. Finally, we conclude in Section VI.
II Tools for exploring non-Markovian strong-coupling quantum thermodynamics
Here, we briefly dive into the key tools for probing quantum thermodynamics in a strongly coupled non-Markovian quantum system. We initiate the discussion from entropy production for a general quantum system surrounded by a bath (using this, we discuss the second law of quantum thermodynamics), followed by the discussion of the change in the system’s and bath’s internal energy and the work done as the mismatch between them. This leads us to the first law of quantum thermodynamics in a general system-bath scenario. In the next step, we discuss how work done is related to the cost of turning the system-bath interactions on or off. To analyze the thermal equilibrium between the system and the bath in the strong-coupling regime, we discuss the formalism of Hamiltonian of mean force and the mean force Gibbs state. Further, we discuss the canonical Hamiltonian of a system.
Throughout the paper, we consider the evolution of the total system SB (system $S$ plus bath $B$) is governed by the Hamiltonian
$H=H_{S}+H_{B}+V,$ (1)
where $H_{S}$, $H_{B}$, and $V$ are the system, bath, and system-bath interaction Hamiltonians, respectively. The dynamics of the total system SB is dictated by the unitary $U=e^{-iHt}$ (for $\hbar=1$) as
$\rho_{SB}(t)=U\left(\rho_{S}(0)\otimes\rho_{B}(0)\right)U^{\dagger},$ (2)
where $\rho_{S}(0)$ and $\rho_{B}(0)$ are the system’s and bath’s initial states, and $\rho_{SB}(t)$ is the joint system-bath state at any time $t$. The reduced state of the system ($\rho_{S}$) or the bath ($\rho_{B}$) is given by
$\rho_{S(B)}={\rm Tr}_{B(S)}\left[U\left(\rho_{S}(0)\otimes\rho_{B}(0)\right)U^{\dagger}\right].$ (3)
II.1 Entropy production, heat and work
Considering the global system-bath dynamics for arbitrary initial states of the system and the bath [Eq. (2)], a formulation for the entropy production was developed in [91, 28]. Here, the dynamical evolution of the system is obtained by tracing out the bath, which is the origin of the irreversibility. The entropy production accounts for the contribution of two processes, one where we discard any information stored locally in the state of the bath and the other where we discard the non-local information shared between the system and the bath, and is given by
Further, when we consider the thermal state to be the initial state of the bath ($\rho_{B}(0)=e^{-\beta H_{B}}/{\rm Tr}\left[e^{-\beta H_{B}}\right]$, where $\beta=1/k_{B}T$ is the inverse temperature of the bath), the entropy production boils down to
$\Sigma=\Delta S_{S}+\beta Q_{B},$ (6)
is the total change in energy of the environment during the unitary evolution of the total system, where $\rho_{B}(t)={\rm Tr}_{S}[\rho_{SB}(t)]$. Equation (6) with $\Sigma\geq 0$ is the standard form of the second law of thermodynamics, and it is in the form of Clausius’ inequality [92]. The derivation of Eq. (6) from Eq. (5) can be done without taking any assumptions about the size of the bath. However, for a finite bath, an expression for the entropy production $\Sigma_{\rm finite}=\Delta S_{S}+\int\delta Q_{B}(t)/T(t)\geq 0$ was suggested in [93, 94], where the temperature $T(t)$ of the finite bath changes with time. Note that if the bath is large enough, such that its temperature stays constant, $\Sigma_{\rm finite}$ reduces to $\Sigma$. Further, although $\Sigma\geq 0$ is still a valid inequality, it was argued that the second law should be identified by $\Sigma_{\rm finite}\geq 0$ [93] for a finite bath. Moreover, $\Sigma_{\rm finite}$ is related to $\Sigma$ by [93]
$\Sigma_{\rm finite}=\Sigma-S\left(\zeta_{B}\left[T(t)\right]||\zeta_{B}\left[T(0)\right]\right),$ (8)
where $\zeta_{B}\left[T(t)\right]=e^{-H_{B}/T(t)}/{\rm Tr}\left[e^{-H_{B}/T(t)}\right]$. In the finite spin bath system studied here, we assume that all bath spins are initially in thermal equilibrium at a temperature $T(0)=T$. As the system evolves, the bath’s temperature changes due to its finite size and its interaction with the central spin.
Next, the change in the internal energy of the system is given by
$\Delta U_{S}=W-Q_{B},$ (10)
[…]
$W$ $={\rm Tr}\left[V\left\{\rho_{SB}(0)-\rho_{SB}(t)\right\}\right].$ (14)
This form is particularly useful as it brings out how, in a general scenario of strong system-bath coupling and under constant Hamiltonians, the work done by the bath can be explicitly computed, saving the need to calculate the change in the internal energies of the system and the bath. It is also worth mentioning that this amount of work must be supplied to turn off (or on) the interaction between the system and the bath.
II.2 The Hamiltonian of mean force
Consider a quantum system $S$ interacting with a thermal bath $B$, which is described by the Hamiltonian $H$ in Eq. (1), The global equilibrium state of the system and the reservoir at temperature $T$ is denoted by ${\zeta}_{SB}$, which is of Gibbs form,
${\zeta}_{SB}:=\frac{\exp(-\beta H)}{{Z}_{SB}},$ (15)
where, ${Z}_{SB}={\rm Tr}\left[\exp(-\beta{H})\right]$ is the partition function for $SB$, and $\beta=(k_{B}T)^{-1}$ ($k_{B}$ is chosen to be one throughout the paper). The interaction term causes the reduced equilibrium state of the system $S$, ${\zeta}^{*}_{S}(T)={\rm Tr}_{B}\left[{\zeta}_{SB}(T)\right]$, to deviate from the thermal state ${\zeta}_{S}(T)=\exp{-\beta H_{S}}/{\rm Tr}\left[\exp{-\beta H_{S}}\right]$ unless the system-bath coupling is very weak. The reduced equilibrium state of the system $S$ as an effective Gibbs state is then given by ${\zeta}^{*}_{S}(T)=\frac{\exp\left[-\beta{\mathcal{H}}_{S}^{*}(T)\right]}{{Z}_{S}^{*}}$, where the partition function for the system S can be expressed as the ratio
${Z}_{S}^{*}={Z}_{SB}/{Z}_{B},$ (16)
where ${Z}_{SB}={\rm Tr}[e^{-\beta{{H}}]}$ and ${Z}_{B}={\rm Tr}_{B}[e^{-\beta{H}_{B}}]$ and,
$\mathcal{{H}}^{*}_{S}(T):=-\frac{1}{\beta}\log\left(\frac{{\rm Tr}_{B}\left[\exp(-\beta{{H}})\right]}{{\rm Tr}_{B}\left[\exp(-\beta{H_{B}})\right]}\right),$ (17)
is the Hamiltonian of mean force [95, 69, 68]. This operator acts as an effective Hamiltonian for the system $S$, incorporating both temperature $T$ and interaction $V$. In the weak coupling regime, it simplifies to the bare Hamiltonian $H_{S}$. The Hamiltonian of mean force (HMF) approach is crucial for understanding key thermodynamic properties like free energy, entropy, and heat capacities in systems with significant environmental interactions [96, 97]. It aids in formulating generalized thermodynamic potentials and investigating equilibrium and non-equilibrium processes, shedding light on quantum dissipation and decoherence. Further, for the thermal equilibrium state $\zeta_{S}(T)$, the entropy of the system is given by $\mathcal{S_{S}}=-{\rm Tr}\left[\zeta_{S}(T)\log\zeta_{S}(T)\right]$ (von Neumann entropy) in the weak coupling regime. In the strong coupling regime due to the alternative partition function ${Z}_{S}^{*}$, the system’s thermodynamic entropy becomes [68]
$\mathcal{S}_{S}=-{\rm Tr}\left[\zeta^{*}_{S}(T)\log\zeta^{*}_{S}(T)\right]+\beta^{2}{\rm Tr}\left[\partial_{\beta}\mathcal{H}_{S}^{*}(T)\zeta_{S}^{*}(T)\right].$ (18)
The first term is the von Neumann entropy of the system, while the second term comes from the temperature dependence of HMF. This suggests that thermodynamic entropy is not universally equivalent to the information content of the equilibrium state in systems with interactions.
II.3 The theory of canonical Hamiltonian
@@
$H^{\rm can}_{S}$ $=\frac{1}{2id}\sum_{j}\gamma_{j}\bigg{(}{\rm Tr}(E_{j})E_{j}^{\dagger}-{\rm Tr}(E_{j}^{\dagger})E_{j}\bigg{)},$ (25)
@@
III The central spin model and its exact dynamics
@@
where $\omega_{0}$ is the transition frequency of the central spin-$1/2$ particle, $\epsilon$ is the interaction strength between the system and the bath with $\sqrt{N}$ being the scaling factor. $\omega/N$ is the scaled frequency of the bath. The evolution of the composite system SB for an arbitrary initial state $\rho_{SB}(0)=\rho_{S}(0)\otimes\rho_{B}(0)$, assuming separable initial condition, of the total system (where $\rho_{S}(0)$ and $\rho_{B}(0)$ are the initial states of the central spin and the bath, respectively) by means of the global unitary $U=e^{-iHt}$ is given by
@@
where $H$ is the total Hamiltonian. Here, we consider the Gibbs state as the initial state of the bath, which is given (using the spectral decomposition of the bath Hamiltonian $H_{B}=\frac{\omega}{N}J_{z}=\frac{\omega}{N}\sum_{n=0}^{N}\left\{\left(\frac{N}{2}-n\right)\ket{n}\bra{n}\right\}=\frac{\omega}{2}\sum_{n=0}^{N}\left\{\left(1-\frac{2n}{N}\right)\ket{n}\bra{n}\right\}$ | }) by
@@
$S_{n}(t)$ $=\sum^{N}_{n=1}e^{-\frac{\beta\omega}{2}\left(1-\frac{2n}{N}\right)}\left\{\left(\frac{\chi_{+}(n)^{2}}{1+\chi_{+}(n)^{2}}\right)^{2}+\left(\frac{\chi_{-}(n)^{2}}{1+\chi_{-}(n)^{2}}\right)^{2}+\left(\frac{\chi_{+}(n)^{2}\chi_{-}(n)^{2}}{\left[1+\chi_{+}(n)^{2}\right]\left[1+\chi_{-}(n)^{2}\right]}\right)2\cos\left(\frac{t\sqrt{b^{2}+4a_{n}^{2}}}{N}\right)\right\}.$ (34)
@@
Figure 1: Variation of the central spin system’s density matrix element $\rho_{00}(t)$ with time for different values of the number of spins $N$ in the spin bath. Subplot (a) has $\epsilon=0.1,T=1$, (b) has $\epsilon=0.5,T=1$, (c) has $\epsilon=0.2,T=10$, and (d) has $\epsilon=0.2,T=0.1$. Further, $\omega_{0}=2.5$ and $\omega=2.0$ for all the plots. The system is initially taken to be in the excited state (|$\ket{\psi(0)}_{S}=\ket{0}_{S}$).
Before exploring the quantum thermodynamic properties of the central spin system, we examine the impact of the number of bath spins $N$, temperature $T$, and coupling strength $\epsilon$ on the system’s dynamics, as can be seen from Fig. 1. We plot the variation in the excited state population $\rho_{00}(t)$ of the system’s state with time for different values of $\epsilon$ and $T$ in the large $N$ limit. $\rho_{00}(t)$ plays an important role in the system’s evolution and is closely connected to the change in its internal energy, see Eq. (41) below. In fact, $\Delta U_{S}(t)=\omega_{0}\left[\rho_{00}(t)-\rho_{00}(0)\right]$, which for initial excited state becomes $\omega_{0}\left[\rho_{00}(t)-1\right]$. From Fig. 1, we observe that as the number of bath spins increases from $N=10^{2}$ to $10^{5}$, approaching the thermodynamic limit, the frequency of oscillations of the excited state population rapidly increases. However, this increase in oscillation frequency is accompanied by a reduction in amplitude over time, particularly for larger $N$. These effects become even more pronounced as the system-bath coupling is increased $\epsilon=0.5$. Additionally, temperature plays a crucial role in influencing the system’s evolution. At low temperatures, the oscillations in the excited state population diminish over time. Notably, in this regime, Fig. 1(d), the system’s steady state is quickly achieved (the oscillations in the excited state population quickly die) as the number of bath spins is increased. This behavior suggests that the coupling strength and the number of bath spins, in conjunction with the bath’s temperature, play a fundamental role in dictating the system’s approach to equilibrium.
IV Exploring the NM quantum thermodynamics of central spin model
Here, we calculate the thermodynamic quantities discussed in Sec. II using the exact dynamics of the central spin model discussed above.
IV.1 Entropy production, heat, work and the laws of quantum thermodynamics
$z(t)$ $=2\left[\alpha(t)\rho_{00}(0)+\eta(t)\rho_{11}(0)\right]-1,$
$x(t)$ $=2\Re\left\{\delta(t)\rho_{01}(0)\right\},$
$y(t)$ $=-2\Im\left\{\delta(t)\rho_{01}(0)\right\}.$ (40)
$\Delta U_{S}=\frac{\omega_{0}}{2}\left[z(t)-z(0)\right].$ (41)
$\rho_{B}(t)={\rm Tr}_{S}\left[e^{-iHt}\left\{\rho_{S}(0)\otimes\rho_{B}(0)\right\}e^{iHt}\right].$ (42)
The variations of the above-mentioned thermodynamic quantities with time are plotted in Fig. 2. We observe that the change in the internal energy of the system $\Delta U_{S}$ is positive and undergoes revivals in time, as depicted by the oscillations. For higher values of system-bath interaction parameter $\epsilon$, we observe well-defined wavepacket-like structures. Further, at a very low value of $\epsilon$, the system’s internal energy change is approximately zero. The above observation matches the variations in the change in the energy of the bath $Q_{B}$ and the work done by the bath $W$. The values of the $Q_{B}$ are approximately zero for weak coupling ($\epsilon/\omega_{0}\approx 0.03$), which is consistent with the GKLS condition, and are otherwise negative, denoting that the energy of the bath decreases during the dynamics, and the positive values of $W$ denote that the bath works on the system. This observation is consistent with the first law of thermodynamics; that is, the energy of the bath decreases while it works on the system. In the process, the internal energy of the system increases. Interestingly, from Fig. 2(c), we observe that the amount of work required to turn off the interaction is approximately zero in the case of weak coupling ($\epsilon/\omega_{0}\approx 0.03$) and increases with an increase in the coupling strength. This observation is consistent with the literature [65].
Figure 3: A comparison of the entropy production $\Sigma$ calculated using Eq. (6) with the entropy production for finite baths $\Sigma_{\rm finite}$ calculated using Eq. (8). The parameters are chosen to be $N=10,T=1,\omega=2,\epsilon=1$, and $\omega_{0}=2.5$.
Also, we calculate the entropy production explicitly for the finite bath $\Sigma_{\rm finite}$ discussed in Eq. (8). To find out the temperature variation with time $T(t)$, we make us of the suggestion in [93] to numerically fit the variable $T$, such that the quantity ${\rm Tr}\left[H_{B}\zeta_{B}\{T(t)\}\right]$ becomes equal to the bath’s internal energy ${\rm Tr}\left[H_{B}\rho_{B}(t)\right]$ at time $t$, Eq. (7). We then use the time-dependent bath temperature $T(t)$ to calculate the quantity $\Sigma_{\rm finite}$. The plot for $\Sigma_{\rm finite}$ is given in Fig. 3. It can be observed that the variation of $\Sigma_{\rm finite}$ is similar to $\Sigma$. The inset plot shows that the entropy production for the finite bath $\Sigma_{\rm finite}$ (red dashed curve) is upper bounded by the quantity $\Sigma$ (teal solid curve), which is consistent with the inequality $\Sigma\geq\Sigma_{\rm finite}$ suggested in [93].
IV.2 HMF and thermodynamic entropy of the central spin model
Here, we use the total Hamiltonian’s spectral decomposition (see Appendix A) to calculate the HMF of the central spin model using Eq. (17). The analytical form of the HMF for the central spin model is given by
$\mathcal{H}_{S}^{*}(\beta)=-\frac{1}{\beta}\begin{pmatrix}\log\left[\pi_{00}(\beta)\right]&0\\ 0&\log\left[\pi_{11}(\beta)\right]\end{pmatrix},$ (43)
where
$\pi_{00}(\beta)$ $=\frac{1}{\xi(\beta)}\left[\sum_{n=1}^{N}\left\{\frac{e^{-\beta\lambda_{\pm}(n)}\chi_{\pm}(n)^{2}}{1+\chi_{\pm}(n)^{2}}\right\}+e^{-\beta(\omega+\omega_{0})/2}\right],$
$\pi_{11}(\beta)$ $=\frac{1}{\xi(\beta)}\left[\sum_{n=1}^{N}\left\{\frac{e^{-\beta\lambda_{\pm}(n)}}{1+\chi_{\pm}(n)^{2}}\right\}+e^{\beta(\omega+\omega_{0})/2}\right],$ (44)
and $\xi(\beta)=\frac{\sinh[\beta\omega(N+1)/2N]}{\sinh(\beta\omega/2N)}$. The forms of the factors $\lambda_{\pm}(n)$ and $\chi_{\pm}(n)$ are given in Appendix A. We further calculate the thermodynamic entropy of the central spin system using the above HMF in Eq. (18). The analytical form of the thermodynamic entropy is given by
$\mathcal{S}_{S}$ $=-\frac{\pi_{00}(\beta)}{\pi_{00}(\beta)+\pi_{11}(\beta)}\log\left(\frac{\pi_{00}(\beta)}{\pi_{00}(\beta)+\pi_{11}(\beta)}\right)$
$-\frac{\pi_{11}(\beta)}{\pi_{00}(\beta)+\pi_{11}(\beta)}\log\left(\frac{\pi_{11}(\beta)}{\pi_{00}(\beta)+\pi_{11}(\beta)}\right)$
$+\beta^{2}\left(\frac{\mu_{00}(\beta)\pi_{00}(\beta)+\mu_{11}(\beta)\pi_{11}(\beta)}{\pi_{00}(\beta)+\pi_{11}(\beta)}\right),$ (45)
where
$\mu_{00}(\beta)$ $=\frac{\log[\pi_{00}(\beta)]}{\beta^{2}}-\frac{\partial_{\beta}\pi_{00}(\beta)}{\beta\pi_{00}(\beta)},$
$\mu_{11}(\beta)$ $=\frac{\log[\pi_{11}(\beta)]}{\beta^{2}}-\frac{\partial_{\beta}\pi_{11}(\beta)}{\beta\pi_{11}(\beta)}.$ (46)
The partial differentiation $\partial_{\beta}\pi_{xx}(\beta)$ for $(x=0,1)$ can be easily calculated from Eq. (44).
Figure 4: Variation of (a) the Hilbert-Schmidt norm of the matrix $\mathcal{H}_{S}^{*}(\beta)-H_{S}$, where $H_{S}$ is the central spin system Hamiltonian and $\mathcal{H}_{S}^{*}(\beta)$ is the HMF, and (b) thermodynamic entropy $\mathcal{S}_{S}$, Eq. (45) as a function of inverse temperature $\beta$. The parameters are: $N=80,\omega=\omega_{0}=5$.
In Fig. 4(a), we show the impact of interaction strength and bath temperature on the HMF (note that here we have taken a large spin bath $N=80$) by plotting the Hilbert-Schmidt norm $\left(\|O\|_{HS}=\sqrt{{\rm Tr}[O^{\dagger}O]}\right)$ of the difference between the system’s bare Hamiltonian $H_{S}$ and HMF $\mathcal{H}_{S}^{*}(\beta)$ as a function of inverse temperature $\beta$ and for different system-bath interaction strengths $\epsilon$. It can be observed from the plots that the HMF becomes equal to the system’s bare Hamiltonian at high temperatures and in the weak coupling regime. The difference is higher for higher $\epsilon$ and lower temperatures. However, the difference saturates as we keep on decreasing the temperature. The thermodynamic entropy of the system is plotted in Fig. 4(b) for different $\epsilon$ as a function of inverse temperature $\beta$. The thermodynamic entropy takes higher values at low temperatures when the system-bath coupling is higher. As we decrease the system-bath coupling, the entropy modulates more slowly for low-temperature values. At very low temperatures, the entropy can be seen to become zero, consistent with the third law of thermodynamics.
IV.3 The canonical Hamiltonian of the central spin model